OpenStudy (anonymous):

Find the 10th term in the sequence. , -1, 2, -4, … a. -256 b. 125 c. 512

OpenStudy (anonymous):

the first term was 1/2 it wouldn't post it

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (solomonzelman):

the common ratio is? the first term is? use: \(\large\color{slate}{\displaystyle a_n=a_1\times r^{n-1}}\)

OpenStudy (solomonzelman):

(or, you can start from -4 ) going like this \(\large\color{slate}{\displaystyle a_{10} = a_3\times r^{10-3}}\)

OpenStudy (solomonzelman):

don't understand my notations ? (if so, please let me know, I will explain them)

OpenStudy (anonymous):

i understand them

OpenStudy (solomonzelman):

My apologies, my notification didn't work.... I didn't know you gave me a medal. I thought you were just sitting there waiting for something to happen. .... you welcome!